# An Investigation of the Current Star Formation Rate of Star-forming Galaxies at z < 0.5 with FADO

### v1

@@ 2.3 Spectral fitting
In this section, we discuss four main issues, which will serve as a foundation for future research in this area. Firstly, we compare our results with the work of Miranda et al. (2023) to explain the significant differences between our main conclusions and the relevant parts of their work. Then, we attempt to identify other factors that could affect SFR estimation. In particular, we also discuss the impact of recent significant findings on the initial mass function on SFR estimation and its trend with redshift. Finally, we look forward to future joint analyses of higher-quality spectroscopic surveys and multi-band spectroscopy.
@@ 4.1 Comparison with Miranda’s results and uncertainty
In comparison to Miranda et al. (2023), who utilized the integrated MPA-JHU as the sample and obtained SFRs with slight or even negligible differences, this survey employed the latest comprehensive galaxy data from SDSS-DR18 and employed the BPT diagram to screen all data, ensuring that all data points are star-forming galaxies (SFGs). This guarantees that all data points exhibit significant differences in nebular contamination. Moreover, the redshift range used in this study, spanning from 0 to 0.5, exceeds the range of 0 to 0.25 used by Miranda et al. (2023). Although in the subsequent fitting process, galaxies with poor spectral quality in the redshift range of 0.4-0.5 were heavily excluded, this range represents the limit of accurately estimating SFR using H6563 in the SDSS survey. In this survey, the redshift range of galaxies is extended to a maximum of 0.5, and preliminary evidence suggests that FADO effectively removes nebular radiation. Additionally, the presence of nebular contamination in SFR calculations increases with increasing redshift. The SFMS also indicates that SFRs derived by FADO align more closely with theoretical predictions. Due to significant errors associated with FADO, we plotted the Figures considering both with and without error scenarios, and the results reveal the anticipated trends under both circumstances. The following section will analyze the causes of large fitting errors.
@@
Both FADO and qsofitmore employ the MC method to obtain uncertainties in emission lines, but there are still some dominant errors inherent in their respective codes. The large errors in FADO arise from its high degrees of freedom, and the error calculation method is relatively simple, involving the superposition of components. FADO fitting includes a greater variety of spectral fitting models and considers a wider range of components. As more statistical uncertainties and degeneracies are introduced among physical components, the comparison between the new code and the old code becomes increasingly complex (Cardoso et al., 2022). The uncertainties in FADO are also derived from random scattering in the MC model, with each step of FADO fitting generating an uncertainty caused by MC. Therefore, for FADO, a comprehensive and complex self-consistent code incorporating many spectral fitting scenarios inevitably generates more uncertainties. When these uncertainties are superimposed, it leads to significant errors observed in the Figures from the previous section. Moreover, there are significant variations in the fitted flux of H6563, with results indicating that removing the flux of H6563 from host galaxies and quasars better fits the trend line of Gaussian model fitting. The closer the variance value of the fitted data to the observed data is to 1, the closer the correlation between the two. Additionally, the number of random scattering points in MC still needs to be considered. FADO iterates the Monte Carlo algorithm more times, while for qsofitmore, we only iterated 15 times. More iterations bring uncertainties closer to reality, but excessive iterations exponentially increase the time taken for qsofitmore to process spectral fitting. To avoid lengthy spectral fitting times, we chose a relatively reasonable value. The greater number of iterations in FADO may be one of the reasons for the larger uncertainty associated with FADO.
@@ 4.2 Other factors affecting SFR estimation
The H6563 line obtained by FADO is generally similar to that of qsofitmore, with a median ratio of 0.09, indicating that the pure stellar fitting model qsofitmore misinterprets a fraction of the nebular emissions as stellar formation activity. Although the overall discrepancy is small, Figure 6 demonstrates a clear trend, preliminarily suggesting an increase in the contribution of nebular emissions with increasing redshift. For the SED model used by FADO, factors such as electron density, extinction, temperature, and nebular kinematics influence the nebular term in the process of fitting spectra with FADO, with electron density and temperature being the dominant factors (Gomes & Papaderos, 2017). Therefore, the underestimation of SFR derived from FADO in this study may be partly attributed to the overestimation of electron density and temperature. Accurate estimation and constraints of galaxy SFR require the combination of multi-band spectral observations due to the distribution of ultraviolet, optical, and infrared radiations in different regions of galaxies and dust obscuration (Elbaz et al., 2018). Active Galactic Nuclei (AGNs) are considered to play a crucial role in shaping and evolving galaxies. The gas on the host galaxy disk interacts with AGN feedback, strongly influencing the overall galaxy SFR. AGNs that may exist in some starburst galaxies can trigger, enhance, or suppress star formation (Elbaz et al., 2018). This also affects the accuracy of estimating galaxy SFR, regardless of the spectral band used.
@@ 4.3 The impact of variable IMF on spectral fitting
Currently, models used for fitting galaxy spectra (e.g., FADO and qsofitmore) assume the initial mass function (IMF) of stars to be constant. However, various indications suggest that the IMF may not be a completely invariant cornerstone of spectral fitting theories. Li et al. (2023) recently announced that the IMF is a function primarily dependent on stellar metallicity rather than a constant as previously assumed. Since stellar metallicity varies among galaxies at different redshifts, the IMF becomes a function of redshift, and thus, future galaxy spectral fitting may need to consider the evolutionary trend of the IMF with redshift. This could profoundly affect the accurate estimation of galaxy SFRs and impose an additional redshift-evolution trend closely related to the IMF on the trend of SFR evolution with redshift.
@@
The impact of IMF on this survey can be described as follows: (a) The luminosity of H6563 is influenced by the IMF. The Lyman continuum (LyC) photon rate is strongly dependent on redshift, and the intensity of nebular emission lines remains a topic of ongoing debate due to inaccurate redshifts. If the Salpeter IMF is used, the luminosity of H6563 will vary by a factor of 4.6 from z = 0.001 to z = 0.040 (Weilbacher & Fritze-v. Alvensleben, 2001). Therefore, calculating SFR using the luminosity of H6563 can lead to significant errors. (b) The influence of metallicity on SFR. Low-metallicity stellar populations are brighter and hotter than solar-metallicity stellar populations with the same IMF and mass constraints. According to the survey by Weilbacher & Fritze-v. Alvensleben (2001), estimating the luminosity of H6563 in low-metallicity dwarf galaxies using empirical calibration may result in an overestimation of the star formation rate by a factor of 3 due to the metallicity effect. (c) Metallicity affects the luminosity of H6563, thereby affecting SFR: a comparison between the IMF of Scalo (1986) and Salpeter (1955) revealed a decrease in H6563 luminosity (Gomes & Papaderos, 2017). (d) The impact of electron density (ne) on SFR. A drawback of all photoionization spectral synthesis (pss) codes is the neglect of ne, which is particularly crucial for modeling pss (Gomes & Papaderos, 2017). Higher electron density implies a higher probability of gas ionization, resulting in more ionized gas. Collisions between them lead to transitions in electron energy levels, ultimately affecting the estimation of SFR.
@@
However, IMF remains an outstanding question in astrophysics to date. The variation in IMF seems not as straightforward as initially imagined. André et al. (2010)’s survey of nearby clouds indicates that the core mass function appears significantly steeper than the cloud mass function and is closer to the mass function of stars. Recently, Li et al. (2023) announced that the IMF is a primary function dependent on stellar metallicity rather than a constant function as previously believed. Since stellar metallicity varies among galaxies at different redshifts, the IMF becomes a function of redshift. Therefore, future galaxy spectral fitting may need to consider the evolutionary trend of the IMF with redshift. Generally, we cannot obtain the most accurate IMF, and we can only reduce the impact of IMF on SFR through other means, which FADO accomplishes reliably. Almost all of these calibrations are based on evolutionary synthesis models, where emerging SEDs are derived for composite stellar populations with specified age combinations, chemical compositions, and IMFs. The calibration by Kennicutt (1998) employs model combinations from the literature and assumes a single power-law IMF with mass limits of 0.1 and 100 (Salpeter, 1955). Compared to a more realistic IMF for H6563, this IMF provides a satisfactory SFR calibration, but for other wavelengths, the relative calibration using different tracers is sensitive to the precise form of the IMF. Therefore, for this survey, we still adopt the IMF form of BC03 and H6563 as the SFR tracer. Different IMFs will affect estimates of star formation rates and stellar mass. However, this effect can be directly addressed by applying a conversion factor between estimates obtained with two IMFs (Zahid et al., 2012).
@@ 4.4 Future spectroscopic data
This investigation ultimately obtained the expected results, demonstrating that SFR estimation of galaxies at higher redshifts is indeed affected by nebular emission. However, this investigation also has a limited sample size for SFR calculations. In addition to expecting improvements in the quality of future SDSS spectral data, supplementary studies using spectral surveys such as LAMOST and DESI have potential. Currently, the S/N of the spectra of the majority of SDSS galaxies with redshifts higher than 0.5 is very low, and the data quality is poor, making it difficult to identify and fit emission lines related to SFR estimation. Therefore, if exploring the SFR of galaxies at higher redshifts, especially those in the cosmic noon period at z 2, not only optical spectra but also UV or IR spectra need to be combined, which will require a huge amount of work. However, there are still doubts about whether FADO has good IR spectral analysis capabilities. We hope that the FADO team can improve and test the code of FADO in the future to perform well not only in IR but also in UV spectra. With the extensive implementation of galaxy IR observation projects by JWST, the era of large sample analysis of high-redshift galaxies will soon arrive. The question of the SFR of the first-generation galaxies in the universe will also be revealed.
@@
In addition, it is imperative to mention that estimating the star formation rate (SFR) of a galaxy through multi-band observations leads to more precise results. The estimation of SFR via H emission is also constrained; H6563 is mostly applicable to galaxies with mean ages ranging from 0.3 to 10 million years (Myr). Comparatively, utilizing 2-10 keV may yield more effective results than H6563. Therefore, for accurate SFR computation, employing multi-band observations is essential to obtain the most comprehensive spectrum of the target galaxy, thereby deriving a complete SFR. Further weight is applied to alternative estimations of attenuation correction and SFR based on UV and IR measurements (Kennicutt & Evans, 2012). Additionally, disregarding the impact of the initial mass function (IMF), significant differences are evident in SFR calculations at H6563 luminosities of (Cerviño et al., 2003). Neglecting the IMF effect for lower SFRs may result in the misestimation of stellar lifetimes, thereby elongating their lifecycle, and inducing larger fluctuations in H emission lines. Hence, employing tracers other than the H emission line to constrain SFR is advisable, such as FUV emission (Kennicutt & Evans, 2012). Observations in the infrared band are crucial for obtaining SFR, as interstellar dust absorbs half of the starlight in galaxies and re-emits it in the infrared. The morphological variations in different dust emission components translate into significant changes in the spectral energy distribution (SED) of dust within and between galaxies (Smith et al., 2007); thus, the conversion of SFR from multi-band observations still requires calibration based on different bands. In terms of correcting SFR after combining optical and infrared observations, Rosario et al. (2016); Kennicutt & Evans (2012) have made notable contributions.
@@
Therefore, to obtain the most accurate SFR for galaxies, it is necessary to combine visible, infrared, and ultraviolet spectra of the target galaxy into a complete galaxy spectrum. However, the objective of this investigation is to compare the effects of two different fitting methods on SFR. Hence, the aforementioned impacts have fairly appeared in both fitting methods, canceling each other out. Nonetheless, it is crucial to note that obtaining accurate SFR requires the combination of correct IMF and multi-band observations.
5 summary
In this tentative work, we demonstrate that FADO, which considers the impact of nebular components on spectral fitting, gives a smaller SFR compared to the pure stellar model of qsofitmore. This further proves the viewpoint of Miranda et al. (2023) that the contribution of nebulae to SFR increases with higher redshift. This may provide some inspiration for later research on higher redshift SFR, where the contribution of nebulae cannot be ignored, especially for SBGs and SFGs. In the future, we expect to use a simultaneous analysis method of IR and optical spectra to obtain more obvious nebular radiation noise. The main conclusions of this study are presented below.
Firstly, about 45 of the data points indicate that qsofitmore underestimated the contribution of nebular emission, resulting in an overestimation of the calculated SFR. Secondly, as redshift increases, the contribution of nebular emission also increases, leading to a further increase in the difference between FADO and qsofitmore.Thirdly, the SFMS diagram shows that FADO’s data has values that are more consistent with theory, further illustrating the advantages of FADO for pure stellar codes. Finally, we predict that at higher redshifts, the difference in nebular emission between FADO and other pure stellar codes will be even greater.
@@ Acknowledgement
The author thank the anonymous referee for their detailed reading of the paper and useful constructive suggestions and comments. This work has made use of Sloan Digital Sky Survey VIII spectroscopic data products. Funding for the Sloan Digital Sky Survey VIII has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-VIII acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS website is www.sdss.org and the spectroscopic data website used in this work is https://dr18.sdss.org/optical/spectrum/search.
@@
The author sincerely thanks Jiahe Xiao for the help with the solution of FADO running problem. Thanks to Ciro Pappalardo and Henrique Miranda from the FADO group for their guidance on the FADO code. Thanks to Yirui Han for supporting the author.
Appendix A Intrinsic Extinction
Extinction generally refers to the process by which photons are absorbed, scattered, and reflected as they pass through a medium, thereby weakening in intensity. In astrophysics, photons from distant galaxies pass through the dust and gas of the galaxy itself, pass through the dust and gas of the Milky Way, and finally reach our telescopes. In order to obtain a more accurate emission line flux and thus calculate the SFR, the extinction of the galaxy itself is necessary. This extinction is called the intrinsic extinction of the galaxy. Follow the Cardelli, Clayton, and Mathis (1988, CCM) extinction definition from (Cardelli et al., 1989):
Where is 3.1, standard value for the diffuse interstellar medium (ISM); represent the extinction at V-band; represent colour excess.
where and are the wavelength-dependent coefficients, which are polynomial of , in unit of (see Eq. 3(a) and 3(b) in Cardelli et al., 1989). is absolute extinction as a function of wavelength. Substitute Equation A2 into Equation A1, and the rearranged equation can be:
Let , the Equation A3 can be written as:
Now, we need to consider the relationship between observed and intrinsic magnitude. Photons from distant galaxies pass through layers of dust and suffer from extinction effects. The intrinsic magnitude should be the difference between the observed magnitude and extinction value:
where represent intrinsic magnitude at wavelength of , while the observed magnitude at wavelength of . Adopting the flux formula, the Equation A5 can be written as:
is the main aim of this section, is the flux after extinction. To achieve this goal, we need to know and . can easily get from spectrum. For can be found via Equation A4. The attenuation of dust on radiation varies in wavelength. Radiation with shorter wavelength suffers heavier attenuation, and vice versa. Therefore, can be calculated by comparing the observed fluxes of two different emission lines. We employ the flux ratio of and to derive the corresponding , which is the Balmer attenuation. Now for and :
Use Equation A7 - Equation A8 we can found:
Substitute Equation A6 into Equation A9, and rearrange equation can get:
and can be found via (Cardelli et al., 1989), ratio of and have been given by FADO and QFM, is 2.86 (Osterbrock, 1989). So, in this way, E(B-V) can be found. Finally, substitute Equation A10 into Equation A6, the can be found:
Example of qsofitmore spectral fitting:

### v3

@@ 3.4 Redshift VS Nebula Ratio
This section discusses four main issues, which will serve as a foundation for future research in this area. Firstly, we compare our results with the work of Miranda et al. (2023) to explain the significant differences between our main conclusions and the relevant parts of their work. Then, we attempt to identify other factors that could affect SFR estimation. In particular, we also discuss the impact of recent significant findings on the initial mass function on SFR estimation and its trend with redshift. Finally, we look forward to future joint analyses of higher-quality spectroscopic surveys and multi-band spectroscopy.
@@ 4.1 Comparison with Miranda’s results and uncertainty
Compared to the results of Miranda et al. (2023), who used an integrated MPA-JHU sample and found negligible differences in SFR, this work leverages the latest SDSS-DR18 comprehensive galaxy data. Through a rigorous selection criterion applied via SQL queries and BPT diagrams, we ensured that all data points were purely star-forming galaxies (SFG). This guarantees that all data points exhibit significant variations in nebular emission contamination. Additionally, the redshift range in this study extends from 0 to 0.4, surpassing the range of 0 to 0.3 used by Miranda et al. (2023), where data beyond a redshift of 0.25 is sparse. Although many galaxy spectra with poor quality in the redshift range of 0.3–0.4 were excluded during the fitting process, this range represents the upper limit for accurately estimating SFR using in the SDSS survey. By extending the redshift range of the galaxies to a maximum of 0.4, preliminary evidence suggests that FADO exhibits less pronounced removal of nebular emission levels at low to intermediate redshifts. Moreover, there are relatively obvious changes in the presence of nebular contamination in the SFR calculations as redshift increases. SFMS further indicates that the stellar mass and SFR obtained with FADO are closer to theoretical predictions. Due to the larger uncertainties of FADO, we observed discrepancies between the two codes in the redshift range of 0–0.1, as shown in Figure 5. However, we attribute these differences to variations in the fitting procedures of the codes rather than the effects of nebular emission contamination. The following section will analyze the reasons for the fitting uncertainties in greater detail.
@@
Both FADO and qsofitmore employ the MC method to obtain uncertainties in emission lines, but some dominant uncertainties are still inherent in their respective codes. The large uncertainties in FADO arise from its high degrees of freedom, and the uncertainties calculation method is relatively simple, involving the superposition of components. FADO fitting includes a greater variety of spectral fitting models and considers a wider range of components. As more statistical uncertainties and degeneracies are introduced among physical components, comparing the new and old codes becomes increasingly complex (Cardoso et al., 2022). The uncertainties in FADO are also derived from random scattering in the MC model, with each step of FADO fitting generating an uncertainty caused by MC. Therefore, for FADO, a comprehensive and complex self-consistent code incorporating many spectral fitting scenarios inevitably generates more uncertainties. When these uncertainties are superimposed, significant uncertainties are observed in the figures from the previous section. Additionally, the number of random scattering points in MC still needs to be considered. FADO iterates the Monte Carlo algorithm more times, while for qsofitmore, we only iterated 15 times. More iterations bring uncertainties closer to reality, but excessive iterations exponentially increase the time taken for qsofitmore to process spectral fitting. To avoid lengthy spectral fitting times, we chose a relatively reasonable value.
@@ 4.2 Other factors affecting SFR estimation
The obtained by FADO is generally similar to that of qsofitmore, with a median ratio of 0.0146 dex. For the SED model used by FADO, factors such as electron density, extinction, temperature, and nebular kinematics influence the nebular term in the process of fitting spectra with FADO, with electron density and temperature being the dominant factors (Gomes & Papaderos, 2017). Therefore, the underestimation of SFR derived from FADO in this study may be partly attributed to the overestimation of electron density and temperature. Accurate estimation and constraints of galaxy SFR require the combination of multi-band spectral observations due to the distribution of ultraviolet, optical, and infrared radiations in different regions of galaxies and dust obscuration (Elbaz et al., 2018). AGNs are considered to play a crucial role in shaping and evolving galaxies. The gas on the host galaxy disk interacts with AGN feedback, strongly influencing the overall galaxy SFR. AGNs that may exist in some starburst galaxies can trigger, enhance, or suppress star formation (Elbaz et al., 2018). This also affects the accuracy of estimating galaxy SFR, regardless of the spectral band used.
@@ 4.3 The impact of variable IMF on spectral fitting
Current galaxy spectral fitting (e.g., FADO and qsofitmore) assumes the initial mass function (IMF) of stars to be constant. However, a recent discovery suggests that the IMF may not be a completely invariant cornerstone of spectral fitting theories. Li et al. (2023) recently announced that the IMF is a function primarily dependent on stellar metallicity rather than a constant as previously assumed. Since stellar metallicity varies among galaxies at different redshifts, the IMF becomes a function of redshift. Thus, future galaxy spectral fitting may need to consider the evolutionary trend of the IMF with redshift. This could profoundly affect the accurate estimation of galaxy SFRs and impose an additional redshift-evolution trend closely related to the IMF on the trend of SFR evolution with redshift.
@@
The impact of the IMF on this survey can be described as follows: (a) The luminosity of is influenced by the IMF. The Lyman continuum (LyC) photon rate is strongly dependent on redshift, and the intensity of nebular emission lines remains a topic of ongoing debate due to inaccurate redshifts. If the Salpeter IMF is used, the luminosity of will vary by a factor of 4.6 from z = 0.001 to z = 0.040 (Weilbacher & Fritze-v. Alvensleben, 2001). Therefore, calculating SFR using the luminosity of can lead to significant uncertainties. (b) The influence of metallicity on SFR. Low-metallicity stellar populations are brighter and hotter than solar-metallicity stellar populations with the same IMF and mass constraints. According to the survey by Weilbacher & Fritze-v. Alvensleben (2001), estimating the luminosity of in low-metallicity dwarf galaxies using empirical calibration may result in an overestimation of the star formation rate by a factor of 3 due to the metallicity effect. (c) Metallicity affects the luminosity of , thereby affecting SFR: a comparison between the IMF of Scalo (1986) and Salpeter (1955) revealed a decrease in luminosity (Gomes & Papaderos, 2017). (d) The impact of electron density (ne) on SFR. A drawback of all photoionization spectral synthesis (pss) codes is neglecting ne, which is particularly crucial for modelling pss (Gomes & Papaderos, 2017). Higher electron density implies a higher probability of gas ionization, resulting in more ionized gas. Collisions between them lead to transitions in electron energy levels, ultimately affecting the estimation of SFR.
@@
However, IMF remains an outstanding question in astrophysics to date. The variation in IMF does not seem as straightforward as initially imagined. André et al. (2010)’s survey of nearby clouds indicates that the core mass function appears significantly steeper than the cloud mass function and is closer to the mass function of stars. Recently, Li et al. (2023) announced that the IMF is a primary function dependent on stellar metallicity rather than a constant function, as previously believed. Since stellar metallicity varies among galaxies at different redshifts, the IMF becomes a function of redshift. Therefore, future galaxy spectral fitting may need to consider the evolutionary trend of the IMF with redshift. The calibration by Kennicutt (1998) employs model combinations from the literature and assumes a single power-law IMF with mass limits of 0.1 and 100 (Salpeter, 1955). Compared to a more realistic IMF for , this IMF provides a satisfactory SFR calibration, but for other wavelengths, the relative calibration using different tracers is sensitive to the precise form of the IMF. Therefore, for this survey, we still adopt the IMF form of BC03 and as the SFR tracer. Different IMFs will affect estimates of star formation rates and stellar mass. However, this effect can be directly addressed by applying a conversion factor between estimates obtained with two IMFs (Zahid et al., 2012).
@@ 4.4 Future spectroscopic data
This work ultimately obtained the expected results; preliminary evidence suggests that estimating the SFR for higher redshift galaxies using is minimally affected by nebular emission contamination. However, this investigation also has a limited sample size for SFR calculations. In addition to expecting improvements in the quality of future SDSS spectral data, supplementary studies using spectral surveys such as DESI have potential. Currently, the S/N of the spectra of the majority of SDSS galaxies with redshifts higher than 0.4 is very low. The data quality is poor, making it difficult to identify and fit emission lines related to SFR estimation. Therefore, if exploring the SFR of galaxies at higher redshifts, especially those in the cosmic noon period at z 2, not only optical spectra but also UV or IR spectra need to be combined, which will require a huge amount of work. However, there are still doubts about whether FADO has good IR spectral analysis capabilities. We hope that the FADO team can improve and test the code of FADO in the future to perform well not only in IR but also in UV spectra. With the extensive implementation of galaxy IR observation projects by JWST, the era of large sample analysis of high-redshift galaxies will soon arrive. The question of the SFR of the first-generation galaxies in the universe will also be revealed.
@@
In addition, it is imperative to mention that estimating the SFR of a galaxy through multi-band observations leads to more precise results. The estimation of SFR via emission is also constrained; is mostly applicable to galaxies with mean ages ranging from 0.3 to 10 million years (Myr). Comparatively, utilizing 2-10 keV may yield more effective results than . Therefore, for accurate SFR computation, employing multi-band observations is essential to obtain the most comprehensive spectrum of the target galaxy, thereby deriving a complete SFR. Further weight is applied to alternative estimations of attenuation correction and SFR based on UV and IR measurements (Kennicutt & Evans, 2012). Additionally, disregarding the impact of the IMF, significant differences are evident in SFR calculations at luminosities of (Cerviño et al., 2003). Neglecting the IMF effect for lower SFRs may result in the misestimation of stellar lifetimes, thereby elongating their lifecycle and inducing larger fluctuations in emission lines. Hence, it is advisable to employ tracers other than the emission line to constrain SFR, such as FUV emission (Kennicutt & Evans, 2012). Observations in the infrared band are crucial for obtaining SFR, as interstellar dust absorbs half of the starlight in galaxies and re-emits it in the infrared. The morphological variations in different dust emission components translate into significant changes in the spectral energy distribution (SED) of dust within and between galaxies (Smith et al., 2007); thus, the conversion of SFR from multi-band observations still requires calibration based on different bands. Regarding correcting SFR after combining optical and infrared observations, Rosario et al. (2016); Kennicutt & Evans (2012) have made notable contributions.
@@
Therefore, to obtain the most accurate SFR for galaxies, combining the target galaxy’s visible, infrared, and ultraviolet spectra into a complete galaxy spectrum is necessary. However, the objective of this investigation is to compare the effects of two different fitting methods on SFR. Hence, the aforementioned impacts have fairly appeared in both fitting methods, cancelling each other out. Nevertheless, it is crucial to note that obtaining accurate SFR requires the combination of correct IMF and multi-band observations.
5 Conclusion
In this preliminary study, we used the latest large-scale SDSS-DR18 survey data, applying stringent criteria to restrict the entire sample to SFG. The redshift range is restricted by the limit for estimating SFR using . Beyond a redshift of 0.4, the quality of emission is poor, and the estimated SFR becomes less accurate. This survey further validates the findings of Miranda et al. (2023), which show that the contribution of nebular emission to the star formation rate increases with redshift. This insight could be valuable for future studies on star formation rates at high redshifts, especially for starburst galaxies and SFGs, where the contribution of nebulae cannot be overlooked. The three main conclusions of this survey are presented below:
(1) The pure stellar code qsofitmore and the FADO code, which accounts for the nebular emission, show consistent results in comparison of the flux. The middle difference is . There is no significant trend in the differences between the two codes concerning the redshift. Therefore, for SFGs with , it seems unnecessary to consider nebular emission in the spectral fitting process. However, for individual source analyses, considering nebular emission results in a reduced flux.
(2) For the trend of nebula ratio with redshift, there is an overall upward trend, particularly in the range of . However, the data points are dispersion, and no clear trend can be concluded. However, the nebula ratio indeed increases with redshift, and more data will be needed to confirm this observation.
(3) The average nebula ratio for the four redshift bins shows a relatively stable upward trend with redshift, and Equation (5) is used to quantify this relationship. We also found that the nebula ratio does not seem to have a strong correlation with stellar mass. For SFGs with , the impact of stellar mass can be somewhat neglected.
Finally, we predict that the difference in nebular emission between FADO and other purely stellar synthesis codes will become even more pronounced at higher redshifts. However, it remains essential to investigate the effect of FADO on the fitting of high-redshift galaxies and the impact of using other emission lines on SFR estimation.
@@ Acknowledgement
The author thanks the anonymous referee for the detailed reading of the paper and useful constructive suggestions and comments. This work has used Sloan Digital Sky Survey VIII spectroscopic data products. The Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions have funded the Sloan Digital Sky Survey VIII. SDSS-VIII acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS website is www.sdss.org, and the spectroscopic data website used in this work is https://dr18.sdss.org/optical/spectrum/search.
@@
The author sincerely thanks Jiahe Xiao for helping with the FADO running problem. The author thanks Han Yirui for all your support.
Appendix A List of SQL query
To find suitable SFG with an obvious velocity offset (velocity offset ), we select galaxies in SDSS DR18 Almeida et al. (2023) with the Structured Query Language (SQL) Search tool (http://skyserver.sdss.org/dr18/en/tools/search/sql.aspx). The applied SQL query in detail is as follows:
SQL Query 1:
SELECT S.plate, S.fiberid, S.mjd,
S.z,G.oiii_5007_flux,
G.h_beta_flux, G.h_beta_flux_err,
G.oiii_5007_flux, G.oiii_5007_flux_err,
G.nii_6584_flux, G.nii_6584_flux_err,
G.h_alpha_flux, G.h_alpha_flux_err,
G.oiii_5007_flux/G.h_beta_flux as o3hb,
G.nii_6584_flux/G.h_alpha_flux as n2ha
FROM GalSpecLine as G
JOIN SpecObjall as S
ON S.specobjid = G.specobjid
WHERE S.class =’GALAXY’ and
S.SNmedian >5 and SNmedian < 20
and S.z <0.2 and z >=0 and S.zwarning = 0 and
S.veldisperr > 0
and S.veldisp > 3*S.veldisperr and
(S.veldisp > 80 and S.veldisp <350)
and G.h_beta_flux_err > 0 and
G.h_beta_flux >5*G.h_beta_flux_err
and G.h_alpha_flux_err > 0 and
G.h_alpha_flux > 5*G.h_alpha_flux_err
and G.oiii_5007_flux_err > 0 and
G.oiii_5007_flux >5*G.oiii_5007_flux_err
and G.nii_6584_flux_err > 0 and
G.nii_6584_flux > 5*G.nii_6584_flux_err
SQL Query 2:
SELECT S.plate, S.fiberid, S.mjd,
S.z, G.oiii_5007_flux,
G.h_beta_flux, G.h_beta_flux_err,
G.oiii_5007_flux, G.oiii_5007_flux_err,
G.nii_6584_flux, G.nii_6584_flux_err,
G.h_alpha_flux, G.h_alpha_flux_err,
G.oiii_5007_flux/G.h_beta_flux as o3hb,
G.nii_6584_flux/G.h_alpha_flux as n2ha
FROM GalSpecLine as G
JOIN SpecObjall as S
ON S.specobjid = G.specobjid
WHERE S.class =’GALAXY’ and
S.SNmedian >5
and S.z <0.4 and z >=0.2 and S.zwarning = 0 and
S.veldisperr > 0
and S.veldisp > 3*S.veldisperr and
(S.veldisp > 80 and S.veldisp <350)
and G.h_beta_flux_err > 0 and
G.h_beta_flux >5*G.h_beta_flux_err
and G.h_alpha_flux_err > 0 and
G.h_alpha_flux > 5*G.h_alpha_flux_err
and G.oiii_5007_flux_err > 0 and
G.oiii_5007_flux >5*G.oiii_5007_flux_err
and G.nii_6584_flux_err > 0 and
G.nii_6584_flux > 5*G.nii_6584_flux_err
Appendix B Example of qsofitmore & FADO spectral fitting
This section shows the comparison between qsofitmore and FADO fitting results for the same galaxy spectrum LEDA 1171303.